# Forecasting and Manipulating the Forecasts of Others

### 12 Mar 2026 v1

@@ 3 Main results
The information wedge $\mathcal{V}^{i}_{t}$ vanishes identically when signals are exogenous to controls, recovering the standard LQG costate equation. When controls enter the state dynamics, a change in signal precision $P^{k}(t)$ also operates through the belief-manipulation channel: it alters how sharply opponents’ posteriors respond to drift perturbations (through $\widetilde{X}^{k}$), which alters the shadow value of shifting those posteriors (through $\bar{H}^{k,i}$), which feeds back into the equilibrium control through (3.3). A model that omits this channel predicts the single-agent response where the true equilibrium response includes a strategic distortion whose mean is given by $\bar{\mathcal{V}}^{i}(t)$ and whose shock-by-shock profile is given by $\mathcal{V}^{i}(t,r)$. The mean wedge $\bar{\mathcal{V}}^{i}(t)$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}(t,r)$ governs the distortion in equilibrium impulse responses. Section 4 quantifies both channels numerically.
@@
If controls do not enter the state dynamics, then $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, the belief-adjoint equations (3.4)–(3.6) decouple, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state.
Lemma 3.4 (Why the maximum-principle stationarity condition is global here).
Fix opponents’ strategy maps $D^{-i}$ in the deterministic-kernel class (hence fixed linear operators from signal histories to actions). Then player $i$’s induced objective $D^{i}\mapsto J^{i}(D^{i};D^{-i})$ is a strictly convex functional on the admissible $L^{2}$ control space. Consequently, the spike-variation stationarity condition (B.8) is not only necessary but sufficient: any admissible control satisfying it is the unique global best response.
@@
With opponents’ maps fixed and linear, the closed-loop drift of $X$ is affine in $D^{i}$, so $(X,D^{-i})$ depend affinely on $D^{i}$. The cost (2.4) is quadratic with $G^{i}_{DD}(t)\succ 0$, hence strictly convex in $D^{i}$. ∎
Corollary 3.5 (Noise-state Volterra best response (global closure)).
Under the conditions of Theorem 3.2, fix an opponents’ profile $D^{-i}$ of noise-state Volterra strategies with deterministic kernels (Definition 2.3). If $D^{i}$ is a best response to $D^{-i}$ over the full admissible $L^{2}$ control class, then for a.e. $t\in[0,T]$,
$G^{i}_{DD}(t)D_{t}^{i}+\bar{H}_{X}^{i}(t)+\int_{0}^{t}H_{X}^{i}(t,u)\,d\widehat{W}_{t}^{i}(u)=0,$
@@
$D_{t}^{i}=-(G^{i}_{DD}(t))^{-1}\bar{H}_{X}^{i}(t)-(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H_{X}^{i}(t,u)\,d\widehat{W}_{t}^{i}(u).$
In particular, every best response over the full admissible class is itself a noise-state Volterra control with deterministic coefficients.
@@
With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [16], Theorem 7.12). Since $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$ (Theorem 3.1 in innovations coordinates), player $i$’s control is a deterministic affine functional of the exogenous process $I^{i}$. The induced objective is strictly convex (Lemma 3.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula. Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the deterministic kernels mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5). ∎
@@
Since every best response over the full admissible $L^{2}$ class is noise-state Volterra (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state Volterra controls is automatically a Nash equilibrium of the unrestricted game. The open question is whether equilibria outside the deterministic-kernel class exist; Section 6 discusses a program for ruling them out.
Remark 3.3 (Single-player benchmark).
When $n=1$, $\mathcal{V}^{i}_{t}\equiv 0$ and the ansatz $H_{X}^{i}(t,r)=S(t)X(t,r)$ reduces the backward system to the standard matrix Riccati equation for $S$, recovering classical separated LQG control.
@@
Replace (2.4) with the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, which penalizes cost variance in addition to cost mean ($\theta_{i}>0$ governs risk aversion; $\theta_{i}\to 0$ recovers (2.4)). The entire Volterra framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large. The filtering closure is unchanged; separation fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel. The full statement and derivation are in Appendix B.6 (Theorem B.1).
@@
$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.8)
where $D^{i,\mathrm{full}}_{t}:=-(G^{i}_{DD}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{i}_{X}(t)+\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
@@
$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{i}_{DD}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.9)
@@
This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{i}_{X}$ inside $\Sigma^{M,i}_{t}$. The joint optimization over $(D^{i},P^{i})$ separates: $D^{i}$ is determined by Corollary 3.5 for any given $P^{i}$, and $P^{i}$ solves
@@
When $n=1$ the FOC for (3.11) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{i}_{X}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
@@
Under time-homogeneous primitives, Volterra kernels reduce to convolution kernels K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals. Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
Reference: kernel objects.
@@
$\widehat{W}_{t}^{i}(u)$ noise-state $X(t,s)$ state impulse response
$F_{t}^{i}(u,s)$ filtering kernel $\widetilde{X}^{i}(t,u)$ unresolved state kernel
$D^{i}(t,u)$ policy kernel $\mathcal{D}^{i}(t,s)$ primitive-noise control
$(\bar{H}_{X}^{i},H_{X}^{i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}(t)+\int_{0}^{t}\mathcal{V}^{i}(t,r)\,dW_{r}$ information wedge
@@ 4 Bilateral belief manipulation in a two-player game
To make the general theory concrete, we specialize to the simplest game that exhibits bilateral belief manipulation: two symmetric players with scalar state, opposing targets, and equal signal precision. This strips away all structure except the core strategic friction. Neither player observes the state directly; each receives a private diffusion signal of precision $p>0$. Because each player’s control enters the state dynamics, it reshapes the opponent’s signal and thereby the opponent’s posterior. The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}(t)$ is the component we visualize below.
@@ 4.2 Numerical equilibrium and kernel visualization
Equilibrium is obtained by relaxed Picard iteration on the deterministic kernel system, discretized on a uniform grid of $N=40$ points; no Monte Carlo is used. At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H_{X}^{i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile. Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{i}_{XX}(T)=0$. The outer residual reaches $10^{-5}$ in $19$ iterations( ${\approx}\,10$ ms in C++). At longer horizons the binding constraint is representational (the filtering kernel becomes more singular) rather than economic.
@@
The key distinction is between $D^{1}(t,s)$ (response to a shock as inferred through the filter).
(a) Primitive-noise state kernel $X(t,s)$.
(b) Equilibrium noise-state feedback kernel $D^{1}(t,s)$.
@@ 4.3 Mean actions, separation failure, and bilateral belief manipulation
Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}(t)$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks. Under perfect information, the mean policy does not depend on signal precision—a separation property. Decentralized information breaks this: changing precision changes how fast opponents’ posteriors react to drift, so players choose actions partly to shift opponents’ beliefs, not only to move the physical state.
Figure 4: Equilibrium mean control $\bar{D}^{1}(t)$ as a function of signal precision $p$, with the perfect-information benchmark (dashed). Low precision makes opponents’ posteriors sluggish, amplifying the incentive to manipulate beliefs; as $p\to\infty$ the mean policy converges to the perfect-information limit.
Deadweight loss.
The kernel figures in Section 4 describe the variance channel (through $\mathrm{Var}(X_{t})$) while Figure 4 isolates the mean channel. In this symmetric game the mean channel captures a purely strategic inefficiency: effort spent shifting the opponent’s posterior raises both tracking loss and effort cost relative to the perfect-information benchmark.
@@
$\bar{\mathcal{V}}^{1}(t)=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}(t,z)\cdot\bar{H}^{2}(t,z)\,dz,$
which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}(t)=-(\bar{X}(t)-1)-\bar{\mathcal{V}}^{1}(t)$. The unresolved kernel $\widetilde{X}^{2}(t,z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}(t,z)$ measures the shadow value to player 1 of that shift. Their inner product $\bar{\mathcal{V}}^{1}(t)$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
Figure 5: Asymmetric equilibrium with $p_{1}=3$ fixed and $p_{2}$ varying. Left: mean controls $\bar{D}^{1}(t)$ (solid) and $\bar{D}^{2}(t)$ (dashed) diverge as the precision gap widens. Center: the mean state path $\bar{X}(t)$ tilts toward the better-informed player’s target. Right: aggregate mean effort $|\bar{D}^{1}|+|\bar{D}^{2}|$ exceeds the perfect-information benchmark (gray dashed), with the excess growing in the precision asymmetry.
Figure 6: Mean information wedges $\bar{\mathcal{V}}^{1}(t)$ and $\bar{\mathcal{V}}^{2}(t)$ as $p_{2}$ varies ($p_{1}=3$ fixed). Both wedges are hump-shaped in $t$, confirming that belief manipulation is most valuable at intermediate dates. Increasing the opponent’s precision amplifies the wedge: a better-informed opponent reacts more sharply to drift, raising the marginal value of manipulating their posterior. The wedges vanish at the terminal date, consistent with the zero terminal condition on the belief adjoint.
Figure 7: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying. Left: competitive targets ($\theta_{1}=1$, $\theta_{2}=-1$). Right: cooperative targets ($\theta_{1}=\theta_{2}=0$). Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}(t)$. Blue: player 1; red: player 2.

### 18 Mar 2026 v2

@@ 3 Main results
The information wedge $\mathcal{V}^{i}_{t}$ vanishes identically when signals are exogenous to controls, recovering the standard LQG costate equation. When controls enter the state dynamics, a change in signal precision $P^{k}(t)$ also operates through the belief-manipulation channel: it alters how sharply opponents’ posteriors respond to drift perturbations (through $\widetilde{X}^{k}$), which alters the shadow value of shifting those posteriors (through $\bar{H}^{k,i}$), which feeds back into the equilibrium control through (3.3). Any model that omits this channel misattributes the strategic distortion to the physical response: the mean error is exactly $\bar{\mathcal{V}}^{i}_{t}$, and the shock-by-shock error is $\mathcal{V}^{i}_{t}(r)$. The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}_{t}(r)$ governs the distortion in equilibrium impulse responses. Section 4 quantifies both channels numerically.
@@
If controls do not enter the state dynamics, then $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, the belief-adjoint equations (3.4) decouple, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state.
Lemma 3.4 (Strict convexity of the best-response problem).
Fix opponents’ strategy maps $D^{-i}$ in the deterministic impulse-response class (hence fixed linear operators from signal histories to actions). Then player $i$’s induced objective $D^{i}\mapsto J^{i}(D^{i};D^{-i})$ is a strictly convex functional on the admissible $L^{2}$ control space. Consequently, the spike-variation stationarity condition (B.8) is not only necessary but sufficient: any admissible control satisfying it is the unique global best response.
@@
With opponents’ maps fixed and linear, the closed-loop drift of $X$ is affine in $D^{i}$, so $(X,D^{-i})$ depend affinely on $D^{i}$. The cost (2.4) is quadratic with $G^{DD,i}(t)\succ 0$, hence strictly convex in $D^{i}$. ∎
Corollary 3.5 (Noise-state linear best response (global closure)).
Under the conditions of Theorem 3.2, fix an opponents’ profile $D^{-i}$ of noise-state linear strategies (Definition 2.3). If $D^{i}$ is a best response to $D^{-i}$ over the full admissible $L^{2}$ control class, then for a.e. $t\in[0,T]$,
$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0,$
@@
$D_{t}^{i}=-(G^{DD,i}(t))^{-1}\bar{H}^{X,i}_{t}-(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u).$
In particular, every best response over the full admissible class is itself a noise-state linear control.
@@
With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [16], Theorem 7.12). Since $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$ (Theorem 3.1 in innovations coordinates), player $i$’s control is a deterministic affine functional of the exogenous process $I^{i}$. The induced objective is strictly convex (Lemma 3.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula. Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the response-maps mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5). ∎
@@
Since every best response over the full admissible $L^{2}$ class is noise-state linear (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state linear controls is automatically a Nash equilibrium of the unrestricted game. The open question is whether equilibria outside the deterministic impulse-response class exist; Section 6 discusses a program for ruling them out.
When $n=1$, $\mathcal{V}^{i}_{t}\equiv 0$ and the ansatz $H^{X,i}_{t}(r)=S(t)X(t,r)$ reduces the backward system to the standard matrix Riccati equation for $S$, recovering classical separated LQG control.
@@
Replace (2.4) with the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, which penalizes cost variance in addition to cost mean ($\theta_{i}>0$ governs risk aversion; $\theta_{i}\to 0$ recovers (2.4)). The entire framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large. The filtering closure is unchanged; certainty equivalence fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel. The full statement and derivation are in Appendix B.6 (Theorem B.1).
@@
$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.6)
where $D^{i,\mathrm{full}}_{t}:=-(G^{DD,i}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
@@
$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{DD,i}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.7)
@@
This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{X,i}$ inside $\Sigma^{M,i}_{t}$. The joint optimization over $(D^{i},P^{i})$ separates: $D^{i}$ is determined by Corollary 3.5 for any given $P^{i}$, and $P^{i}$ solves
@@
When $n=1$ the FOC for (3.9) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{X,i}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
@@
Under time-homogeneous primitives, response maps become time-invariant K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals. Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
Reference: impulse-response objects.
@@
$\widehat{W}_{t}^{i}(u)$ noise-state $X_{t}(s)$ state impulse response
$F_{t}^{i}(u,s)$ filtering impulse response $\widetilde{X}^{i}_{t}(u)$ unresolved state response
$D^{i}_{t}(u)$ policy impulse response $\mathcal{D}^{i}_{t}(s)$ primitive-noise control
$(\bar{H}^{X,i},H^{X,i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r}$ information wedge
@@ 4 Bilateral belief manipulation in a two-player game
To make the general theory concrete, we specialize to the simplest game that exhibits bilateral belief manipulation: two symmetric players with scalar state, opposing targets, and equal signal precision. This strips away all structure except the core strategic friction. Neither player observes the state directly; each receives a private diffusion signal of precision $p>0$. Because each player’s control enters the state dynamics, it reshapes the opponent’s signal and thereby the opponent’s posterior. The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}_{t}$ is the component we visualize below.
@@ 4.2 Numerical equilibrium and kernel visualization
Equilibrium is computed by iterating the best-response operator on the deterministic impulse-response system, discretized on a uniform grid of $N=40$ points; no Monte Carlo simulation is needed. At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H^{X,i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile. Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{XX,i}(T)=0$. The outer residual reaches $10^{-5}$ in $19$ iterations( ${\approx}\,10$ ms in C++). At longer horizons the binding constraint is representational (the filtering kernel becomes more singular) rather than economic.
@@
The key distinction is between $D^{1}_{t}(s)$ (response to a shock as inferred through the filter).
(a) Primitive-noise state kernel $X_{t}(s)$.
(b) Equilibrium noise-state feedback kernel $D^{1}_{t}(s)$.
@@ 4.3 Mean actions, separation failure, and bilateral belief manipulation
Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}_{t}$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks. Under perfect information, the mean policy does not depend on signal precision—a certainty-equivalence property. Decentralized information breaks this: changing precision changes the speed at which opponents update, giving each player a reason to act on posteriors rather than just on the physical state.
Figure 4: Equilibrium mean control $\bar{D}^{1}_{t}$ as a function of signal precision $p$, with the perfect-information benchmark (dashed). Low precision makes opponents’ posteriors sluggish, amplifying the incentive to manipulate beliefs; as $p\to\infty$ the mean policy converges to the perfect-information limit.
@@
$\bar{\mathcal{V}}^{1}_{t}=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}_{t}(z)\cdot\bar{H}^{2}_{t}(z)\,dz,$
which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}_{t}=-(\bar{X}_{t}-1)-\bar{\mathcal{V}}^{1}_{t}$. The unresolved kernel $\widetilde{X}^{2}_{t}(z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}_{t}(z)$ measures the shadow value to player 1 of that shift. Their inner product $\bar{\mathcal{V}}^{1}_{t}$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
Figure 5: Asymmetric equilibrium with $p_{1}=3$ fixed and $p_{2}$ varying. Left: mean controls $\bar{D}^{1}_{t}$ (solid) and $\bar{D}^{2}_{t}$ (dashed) diverge as the precision gap widens. Center: the mean state path $\bar{X}_{t}$ tilts toward the better-informed player’s target. Right: aggregate mean effort $|\bar{D}^{1}|+|\bar{D}^{2}|$ exceeds the perfect-information benchmark (gray dashed), with the excess growing in the precision asymmetry.
Figure 6: Mean information wedges $\bar{\mathcal{V}}^{1}_{t}$ and $\bar{\mathcal{V}}^{2}_{t}$ as $p_{2}$ varies ($p_{1}=3$ fixed). Both wedges are hump-shaped in $t$, confirming that belief manipulation is most valuable at intermediate dates. Increasing the opponent’s precision amplifies the wedge: a better-informed opponent reacts more sharply to drift, raising the marginal value of manipulating their posterior. The wedges vanish at the terminal date, consistent with the zero terminal condition on the belief adjoint.
Figure 7: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying. Left: competitive targets ($\theta_{1}=1$, $\theta_{2}=-1$). Right: cooperative targets ($\theta_{1}=\theta_{2}=0$). Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}_{t}$. Blue: player 1; red: player 2.